A certain gas expands at a constant pressure of 2.7×10^5 Pa. If the work done by the system is 600 J, what is the change in the gas volume? Express your answer with the appropriate units.
Answer #1

Work done dW=Pdv

given dW = 600 J, P = 2.7x10^5 Pa

dV = dW/P = 600/2.7x10^5 = 222.2x10^-5

or 2.22X10^-7 m^3